OpenStudy (anonymous):

Fourth root of 9 * square root of 9 / fourth root of 9 to the power of 5

OpenStudy (anonymous):

A. 5 to the power of negative 1 over 6 B. 5 to the power of 3 over 2 C. 9 D. 92

OpenStudy (anonymous):

@Destinymasha @fluteprincess352 @Conqueror @sammixboo @Angel_Bear19 @winsteria

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

@k_lynn @sammixboo @iambatman

OpenStudy (anonymous):

@GreenCat @hhelpplzzzz @ImaGuest @jhonyy9 @x3_x3 @mg_omg @Nnesha @bohotness @YanaSidlinskiy @rayman5766 @pooja195 @Hotchellerae21 @uri

jigglypuff314 (jigglypuff314):

could you try using the blue "Equation" button to make it more clear, I am not getting any of the options given ^_^

OpenStudy (anonymous):

Please!

jigglypuff314 (jigglypuff314):

\[\large \frac{ \sqrt[4]{9}*\sqrt{9} }{ (\sqrt[4]{9})^5 }\]is this what you meant?

OpenStudy (anonymous):

5 is inside the perenthesis

OpenStudy (anonymous):

The rest is good

OpenStudy (anonymous):

@jigglypuff314

OpenStudy (anonymous):

@fani1996 @jojo4eva @confluxepic

OpenStudy (anonymous):

@jinxhead20

OpenStudy (anonymous):

pellet! I came late ....